A 30-foot board is sawed into two pieces.One piece is 6 feet longer that the other. How long are the two pieces?
Answer by checkley71(8403)

x+(x+6)=30 or 2x+6=30 or 2x=24 or x=24/2 or x=12 for the short board & 12=6=18 for the longer board.
proof 12+18=30 or 30=30